# Evaluating and Advancing Multimodal Large Language Models in Ability Lens

### v1

@@ Abstract
As multimodal large language models (MLLMs) advance rapidly, rigorous evaluation has become essential, providing further guidance for their development. In this work, we focus on a unified and robust evaluation of vision perception abilities, the foundational skill of MLLMs. We find that existing perception benchmarks, each focusing on different question types, domains, and evaluation metrics, introduce significant evaluation variance, complicating comprehensive assessments of perception abilities when relying on any single benchmark. To address this, we introduce AbilityLens, a unified benchmark designed to evaluate MLLMs across six key perception abilities, focusing on both accuracy and stability, with each ability encompassing diverse question types, domains, and metrics. With the assistance of AbilityLens, we: (1) identify the strengths and weaknesses of current models, highlighting stability patterns and revealing a notable performance gap between open-source and closed-source models; (2) introduce an online evaluation mode, which uncovers interesting ability conflict and early convergence phenomena during MLLM training; and (3) design a simple ability-specific model merging method that combines the best ability checkpoint from early training stages, effectively mitigating performance decline due to ability conflict. The benchmark and online leaderboard will be released soon.
footnotetext: Work done by Feng Chen during his internship at TikTok, Australia.
@@ 1 Introduction
(a) Evaluation variance across benchmarks on grounding capability.
(b) Comparison on grounding capability using our AbilityLens.
Figure 1: Motivation behind our AbilityLens and our solution. (a) shows evaluation variance across five existing benchmarks [12, 19, 48, 9, 42] on grounding capability. The differing best and worst models across benchmarks underscore the need for a holistic assessment of perception abilities. (b) Our AbilityLens offers a more complete evaluation by encompassing diverse question types, domains, and metrics, assessing both accuracy and stability of MLLMs.
@@
The success of large language models (LLM) [41, 6, 8, 2] has inspired strong interest in developing multimodal large language models (MLLMs) [20, 23, 11, 43], equipping it with emergent abilities to tackle various vision-language tasks [31, 30, 28]. With the rapid development of these models, rigorous evaluation [12, 49, 48] has become crucial, as it highlights existing shortcomings and guides further model improvement. For vision-language cross-modality, evaluating vision perception is particularly critical [40], as it establishes the foundation to understand how effectively MLLMs interpret and comprehend the visual world.
@@
Recent benchmarks [12, 9, 31] evaluate perception capabilities by focusing on varied question types, domains, and metrics, aiming to capture multiple facets of model performance. However, these differences inevitably introduce evaluation variance. As shown in Fig. 1 (a), benchmarks like MME [12], MMBench [48], and MuirBench [42] yield distinct best and worst models, making it challenging to comprehensively assess perception capabilities with any single benchmark. Additionally, unlike conventional assessments that overemphasize model accuracy [17, 52], we argue that stability [5, 47], achieving consistent performance across diverse factors such as domains, question types, and metrics, is also important. For example, a robust model excelling in answering True/False questions in MME [12] should also handle well multiple-choice questions in SeedBench [19]. Thus, current benchmarks are still insufficient for a comprehensive evaluation of perception capability.
@@
In this paper, we introduce AbilityLens, a broad-range and efficient benchmark designed to solve these issues. Building on prior perception categories [12, 42, 9], AbilityLens assesses six key perception abilities: counting, OCR, attribute recognition, entity extraction, grounding, and structural data understanding. Our benchmark is constructed through a systematic process. First, to ensure adequate data for each question type, metric, and domain, we compile data from 11 existing benchmarks [12, 48, 19, 42, 9, 25, 18, 30, 17, 52], resulting in over 1,000 test samples per ability type and a total of 12,000 test samples overall. Second, we derive sub-metrics from source benchmarks and unify them into a single composite score to measure accuracy and stability. Recognizing that different question types bring distinct baseline performance, e.g., True/False questions have a natural accuracy of 50% with random choice while four-choice questions only have 25%, we apply baseline correction to each sub-metric to align performance levels. Then, model accuracy is calculated by a weighted sum of these sub-metrics while model stability can be measured via the variance of them. To this end, our AbilityLens provides a more comprehensive view on perception capabilities over previous benchmarks, as shown in Fig. 1 (b).
In addition, our AbilityLens provides both online and offline evaluation modes to accommodate diverse model assessment needs. Online evaluation mode aims to monitor the training dynamic by evaluating the intermediate saved training checkpoints, so as to adjust the training strategy to advance MLLMs. Different from current benchmarks [12, 48], our AbilityLens provides a more granular understanding of how abilities evolve and interact with each other. Offline evaluation mode is to compare MLLMs in each perception ability via absolute ranking. We report 14 state-of-the-art MLLMs [20, 43, 10, 13, 4] on AbilityLens and examine the benefits brought by LLM size, series, and training data. We summarize the key experimental findings on AbilityLens: (1) Although current state-of-the-art MLLMs like LLaVA-OV-72b [20] and Qwen2-VL-72b [43] outperform commercial models [13, 4] in accuracy, there is a notable stability gap between them, emphasizing the importance of evaluating and improving stability alongside accuracy. (2) We observe the phenomenon of early convergence and potential conflict in ability. Specifically, different perception abilities exhibit different improvement curves, with some abilities even decreasing in both accuracy and stability after further training, which is considered an ability conflict. Additionally, inspired by the model soup [44], we conduct preliminary research on leveraging early convergence to mitigate the performance decline caused by conflict. We propose a simple strategy that combines the final model with checkpoints excelling in target abilities at early training stages, achieving better results than continuous training with minimal computation cost.
In summary, we highlight the difference with current benchmarks in  Tab. 1 and list our contributions as:
(1) We introduce AbilityLens, a comprehensive visual perception benchmark that identifies the strengths and weaknesses of MLLMs in both accuracy and stability.
(2) We offer both online and offline evaluation modes to monitor training dynamics and provide fine-grained rankings of state-of-the-art MLLMs.
(3) We propose a simple yet effective ability-specific model merging method to mitigate the observed ability conflict with minimal additional cost.
Figure 2: Overview of AbilityLens: An ability-focused benchmark that (a) provides comprehensive ranking of VLMs across six perception capabilities through offline evaluation, (b) tracks training dynamics to identify optimization trends and potential conflict via online monitoring, and (c) introduces an efficient ability-specific model merging method to enhance targeted capabilities while preserving others.
@@ 2 Related Work
Ability-perspective MLLM benchmark. To enable a systematic evaluation of MLLM advancements, recent benchmarks [33, 3] aim to expose granular weaknesses within specific capabilities by regrouping benchmarks [21, 31, 49] according to model abilities. For instance, Unibench [3] categorizes 53 vision-language benchmarks by different capabilities, though most focus on traditional recognition tasks and are limited in evaluating state-of-the-art MLLMs. EUREKA [7], on the other hand, concentrates on non-saturated samples within each ability, offering a more targeted view of areas most ripe for improvement. However, these current benchmarks primarily emphasize offline model comparisons, providing limited insight into when and why models may underperform on specific capabilities or efficient ways to enhance these weaker perception capabilities.
Model merging and model soup. Model merging [46, 38] aims to consolidate multiple parameter sets into a single model without retraining, often serving as a method for continual learning. For instance, model soup [44] proposes to greed search two fine tuned models for averaging weights to achieve a better model. In LLMs, Task Arithmetic [14] introduces task vectors to modify pre-trained models through arithmetic operations on weight differences, improving task performance without retraining. Building on these ideas, we present a post-processing, ability-specific model merging pipeline designed to enhance weaker perception abilities in MLLMs. This method effectively alleviates ability conflict, minimizing the need for extensive dataset modifications by leveraging the saved checkpoints during model development.
@@ 3 AbilityLens
Table 2: The data component detail. Data of AbilityLens is sourced from 11 public benchmarks. We reduce the sample size in large datasets, where $\rightarrow$ indicates the number of samples before and after reduction.
In this section, we elaborate on the details of our AbilityLens. As shown in Fig. 2, our AbilityLens includes (1) dataset collection and metric, (2) evaluation framework with online and offline modes, and (3) ability-specific model merging to improve weak abilities.
@@ 3.1 Dataset Collection and Metrics
Dataset collection. A major challenge lying in building a comprehensive ability-focused evaluation is the lack of a unified evaluation protocol, as model and benchmark creators often implement custom pipelines that vary widely in data preparation, output postprocessing, and metric calculation. Therefore, we develop our AbilityLens based on LMMs-Eval [51], a comprehensive evaluation suite for VLMs that supports over 90 VLM benchmarks. In detail, as illustrated in Table 2, AbilityLens focuses on 6 core perception abilities following prior works [12, 48]: counting, OCR, attribute recognition, entity extraction, grounding, and structural data understanding. We select relevant 11 datasets from LMMs-Eval to construct a uniform one for each ability. The included benchmarks encompass MME [12], MMBench [48], SeedBench [19], MuirBench [42], Mirb [53], MMStar [9], OCRBench [25], SynthDog [18], ChartQA [30], AI2D [17], and ConBench [52], covering diverse domain, question types and metrics. To achieve efficient yet broad-ranging evaluation, we follow LMMs-Eval [51] to remove the redundant samples, allowing users to obtain the results of a 7B MLLM in 0.5 hours with 8 H100 GPUs.
@@
Evaluation metric. For each ability, we introduce two metrics: one for model accuracy and another for modal stability, both derived from sub-metrics of the source benchmarks. Because different question types have varying baseline performance levels, representing the expected accuracy from random guessing, we first adjust each sub-metric to account for these differences. For instance, True/False questions have a baseline accuracy of 50% due to the binary choice, while multiple-choice questions may have a lower baseline depending on the number of options. To ensure that the true ability score remains independent of question type, we apply baseline correction to each sub-metric, aligning them to a consistent standard. Let $m_{i}$ represent the value of the $i$-th sub-metric, with $\text{UB}_{i}$ denoting its upper bound and $\text{BL}_{i}$ its baseline performanceThe baseline performance for the VQA benchmark is set to 0.. The baseline correction BC for metric $m_{i}$ is then defined as:
$\text{BC}(m_{i})=\frac{m_{i}-\text{BL}_{i}}{\text{UB}_{i}-\text{BL}_{i}}.$ (1)
@@
The overall accuracy, denoted as $\mathcal{A}$, is then calculated as a weighted sum of sub-metrics, weighted by their sample counts:
$\mathcal{A}=\sum_{i}\frac{n_{i}\cdot m_{i}}{N},$ (2)
@@
where $N$ is the total sample count and $n_{i}$ is the sample count for metric $m_{i}$. To assess model stability across each ability, we compute the stability score $\mathcal{I}$ as the standard deviation of the sub-metrics, with higher values indicating greater instability:
$\mathcal{I}=\text{std}(m).$ (3)
@@ 3.2 Evaluation Framework
Figure 3: Monitoring training dynamics with AbilityLens. Unlike traditional methods that focus solely on average scores, often obscuring underlying conflicts, AbilityLens tracks optimization trends for each specific ability, revealing an ability conflict indicated by a significant drop in accuracy and stability of OCR performance after 10,000 steps.
@@
Offline Evaluation Mode provides a comprehensive assessment of each model’s perception capabilities in accuracy and stability, as shown in Tab. 3 and Tab. 4, respectively. We evaluate state-of-the-art MLLMs with various model sizes and training data, including both commercial and open-source models, across six core abilities.
Online Evaluation Mode enables real-time monitoring of training dynamics throughout model development, leveraging AbilityLens for a more unified and comprehensive assessment compared to existing benchmarks. By evaluating checkpoints at various training stages, we can (1) assess the impact of training configuration adjustments, (2) identify optimal checkpoints for specific abilities, (3) track trends in ability-specific optimization, and (4) detect potential conflicts between abilities. In Fig. 3, we illustrate AbilityLens in action through a reimplementation of LLaVA-OV-SI’s [20] single-stage training process. Traditional approaches, represented by the fourth subfigure in Fig. 3, focus solely on average scores and often obscure underlying conflicts. In contrast, AbilityLens offers a capability-decomposed analysis tool that enables researchers to gain deeper insights and make targeted improvements in their VLMs. For example, after 10,000 training steps, the OCR ability displays a marked drop in both accuracy and stability, revealing an ability conflict that conventional methods might overlook. AbilityLens thus not only uncovers such critical issues but also supports targeted adjustments, as we demonstrate with an efficient resolution in the following section.
3.3 Ability-specific Model Merging
While continuous fine-tuning [16] on relevant datasets is a straightforward approach to strengthening weaker abilities in MLLMs, this method requires substantial computational resources and presents challenges in optimizing training configurations, such as data mixing ratios and learning rates. To address these limitations, we introduce Ability-specific Model Merging (ASMM), a straightforward yet effective strategy that merges checkpoints excelling in target abilities at earlier training stages. ASMM is designed to enhance weak capabilities by inheriting knowledge from high-performing checkpoints specific to these abilities, while simultaneously preserving overall model performance and significantly reducing computational costs.
After identifying ability conflicts with AbilityLens, as shown in Fig. 3, we designate the final step checkpoint as the base model $\theta_{\text{base}}$, as it generally achieves strong performance across most abilities. We then select a high-performing checkpoint $\theta_{\text{checkpoint}}$ for the target ability from an earlier training stage and merge these two models into a unified model via linear interpolation:
$\theta_{\text{interpolate}}=\alpha\cdot\theta_{\text{base}}+(1-\alpha)\cdot\theta_{\text{checkpoint}},$ (4)
where $\alpha$ is the merging ratio and $\theta_{\text{interpolate}}$ is the final model after merging. To determine the optimal value of $\alpha$, we conduct a grid search over the range $[0.1,0.9]$ with a step size of 0.1. As each AbilityLens evaluation requires only 0.5 hours per iteration, the search can be completed within approximately 5 hours using 8 H100 GPUs.
Table 3: Accuracy comparison. Accuracy $\uparrow$ of 14 MLLMs on AbilityLens where cells highlighted in green indicate the best performance, and red indicate the worst performance.
@@ 3.3 Ability-specific Model Merging
QwenVL2-2b 48.28 73.36 35.00 82.39 56.61 46.36 57.00
QwenVL2-7b 50.95 79.29 47.12 85.10 62.66 67.56 65.45
QwenVL2-72b 55.84 86.09 58.70 86.43 65.11 79.32 71.92
@@
Table 4: Stability comparison. Stability score $\downarrow$ of 14 MLLMs on AbilityLens, with cells highlighted in red for the highest and green for the lowest scores. Lower scores indicate better stability.
@@
LLaVA1.5-7b 26.43 11.91 16.32 12.42 21.58 20.43 18.18
LLaVA1.6-7b 25.36 24.35 14.87 11.66 23.86 22.74 20.47
LLaVA-OV-0.5b 21.09 47.36 13.90 25.39 24.59 22.77 25.85
LLaVA-OV-7b 26.95 30.71 18.08 9.99 24.63 8.99 19.89
LLaVA-OV-SI-7b 33.64 10.86 25.56 6.92 27.01 12.26 19.38
LLaVA-Video-7b 16.70 39.13 10.46 44.80 22.54 17.27 25.15
LLaVA-OV-72b 16.84 18.57 19.56 8.50 19.86 12.22 15.92
QwenVL2-2b 16.91 40.89 19.56 14.79 29.67 18.78 23.43
QwenVL2-7b 18.95 11.81 19.14 13.07 25.66 13.24 16.98
QwenVL2-72b 23.13 10.72 16.17 11.72 24.13 3.05 14.82
InternVL2-8b 20.23 16.13 17.04 11.83 27.15 11.22 17.27
gpt-4-vision-preview 26.20 11.52 12.60 27.96 13.03 11.52 17.14
gpt-4o-2024-08-06 20.66 7.85 6.39 12.12 9.99 14.07 11.85
claude-3-5-sonnet-2024102 18.02 9.92 5.88 12.25 11.78 12.58 11.74

### v2

# Evaluating and Advancing Multimodal Large Language Models in Perception Ability Lens

@@ Abstract
As multimodal large language models (MLLMs) advance rapidly, rigorous evaluation has become essential, providing further guidance for their development. In this work, we focus on a unified and robust evaluation of vision perception abilities, the foundational skill of MLLMs. We find that existing perception benchmarks, each focusing on different question types, domains, and evaluation metrics, introduce significant evaluation variance, complicating comprehensive assessments of perception abilities when relying on any single benchmark. To address this, we introduce AbilityLens, a unified benchmark designed to evaluate MLLMs in six key perception abilities (ranging from counting, OCR, to understanding structural data), focusing on both accuracy and stability, with each ability encompassing diverse types of questions, domains, and metrics. With the assistance of AbilityLens, we: (1) identify the strengths and weaknesses of current main-stream MLLMs, highlighting stability patterns and revealing a notable performance gap between state-of-the-art open-source and closed-source models; (2) uncover interesting ability conflict and early convergence phenomena during MLLM training; (3) reveal the primary reason of ability conflict is data mixing ratio and LLM model size; and (4) discuss the effectiveness of some straightforward strategies e.g., fine-tuning and model merging, to solve the ability conflict. The benchmark and online leaderboard is released in https://github.com/Chenfeng1271/AbilityLens.
(a) Evaluation variance across three benchmarks on entity capability.
(b) Comparison on entity capability using our AbilityLens.
Figure 1: Motivation behind our AbilityLens and our solution. (a) shows evaluation variance across three existing benchmarks [12, 18, 40] on entity capability. The differing best and worst models across benchmarks underscore the need for a holistic assessment of perception abilities. (b) Our AbilityLens offers a more complete evaluation by encompassing diverse question types, domains, and metrics, assessing both accuracy and stability of MLLMs. The sizes of closed-source models are artificial estimates.
@@ 1 Introduction
The success of large language models (LLM) [36, 5, 8, 1] has inspired strong interest in developing multimodal large language models (MLLMs) [19, 23, 11, 38], equipping it with emergent abilities to tackle various vision-language tasks [29, 28, 26]. The key capability that distinguishes MLLMs from LLMs is their ability to ‘see’, i.e., their perception skills. Thus, evaluating vision perception is particularly critical [35], as it establishes the foundation to understand how effectively MLLMs interpret and comprehend the visual world.
@@
Recent benchmarks [12, 9, 29] evaluate perception capabilities by focusing on varied question types, domains, and metrics, aiming to capture multiple facets of model performance. However, these differences inevitably introduce evaluation variance. As shown in Figure 1 (a), benchmarks like MME [12], MMBench [40], and SeedBench [18] yield distinct best and worst models, making it challenging to comprehensively assess perception capabilities with any single benchmark. Therefore, unlike conventional assessments that overemphasize model accuracy [16, 44], we argue that stability [4, 39], achieving consistent performance across diverse factors such as domains, question types, and metrics, is also important. This metric quantifies the evaluation variance exhibited by models across diverse evaluation scenarios. For example, a robust model excelling in answering True/False questions in MME [12] should also handle well multiple-choice questions in SeedBench [18]. Thus, current benchmarks are still insufficient for a comprehensive evaluation of perception capability.
In this paper, we introduce AbilityLens, a broad-range and efficient benchmark designed to solve these issues. Building on prior perception categories [12, 37, 9], AbilityLens assesses six key perception abilities: counting, OCR, attribute recognition, entity extraction, grounding, and structural data understanding. Our benchmark is constructed through a systematic process. First, to ensure adequate data for each question type, metric, and domain, we compile data from 11 existing benchmarks [12, 40, 18, 37, 9, 24, 17, 28, 16, 44], resulting in over 1,000 test samples per ability type and a total of 12,000 test samples overall. Second, we derive sub-metrics from source benchmarks and unify them into a single composite score to measure accuracy and stability. Recognizing that different question types bring distinct baseline performance, e.g., True/False questions have a natural accuracy of 50% with random choice while four-choice questions only have 25%, we apply baseline correction to each sub-metric to align performance levels. Then, the model accuracy of each ability is calculated by a weighted sum of these corresponding corrected sub-metrics. Furthermore, we assess the stability of the target model by computing the variance of its z-scores across sub-metrics, which directly reflects its relative performance compared to all candidate models on each sub-metric. To this end, our AbilityLens provides a more comprehensive view on perception capabilities over previous benchmarks, as shown in Figure 1 (b) and Table 1.
Considering our AbilityLens is more general than existing benchmarks, we further use it to monitor the training dynamic by evaluating the intermediate saved training checkpoints, so as to adjust the training strategy to advance MLLMs. Different from current benchmarks [12, 40], our AbilityLens provides a more granular understanding of how abilities evolve and interact with each other. Besides, we also compare mainstream MLLMs in each perception ability via absolute ranking. We report 18 state-of-the-art MLLMs [19, 38, 10, 13, 3] on AbilityLens and examine the benefits brought by LLM size, series, and training data. We summarize the key experimental findings on AbilityLens: (1) Although current state-of-the-art MLLMs like LLaVA-OV-72b [19] and Qwen2VL-72b [38] outperform commercial models [13, 3] in accuracy, there is a notable stability gap between them as shown in Figure 1 (b), emphasizing the importance of evaluating and improving stability alongside accuracy. (2) We observe the phenomenon of early convergence and potential conflict in abilities during model training. Specifically, different perception abilities exhibit different improvement curves, with some abilities meeting performance decline in both accuracy and stability after further training, which is considered an ability conflict. (3) After extensive experiments across various combinations of visual encoders, LLMs, and datasets, we conclude that the data mixing ratio and LLM model size are the primary factors behind the observed ability conflict. Additionally, we conduct a preliminary discussion on leveraging existing fine-tuning and model merging methods to solve the ability conflict.
@@ 2 Related Work
Perception-based MLLM benchmarks. In parallel with architectural advancements [19, 6], considerable effort has been dedicated to improving benchmarks for MLLMs, which play a crucial role in steering the development of the next generation of models [41]. As a fundamental capability, perception is essential to the performance of MLLMs. To evaluate this skill, benchmarks such as MME [12], MMStar [9], MMBench [40], and SeedBench [18] have been introduced, aiming to capture data from diverse domains and cover a wide range of question types. However, these differences inevitably introduce evaluation variance, motivating the development of a unified dataset that enables a comprehensive assessment of perception abilities.
Ability-perspective MLLM benchmarks. To enable a systematic evaluation of MLLM advancements, recent benchmarks [30, 2] aim to expose granular weaknesses within specific capabilities by regrouping benchmarks [20, 29, 41] according to model abilities. For instance, Unibench [2] categorizes 53 vision-language benchmarks by different capabilities, though most focus on traditional recognition tasks and are limited in evaluating state-of-the-art MLLMs. EUREKA [7], on the other hand, concentrates on non-saturated samples within each ability, offering a more targeted view of areas most ripe for improvement. However, these current benchmarks primarily emphasize offline model comparisons, providing limited insight into when and why models may underperform on specific capabilities or efficient ways to enhance these weaker perception capabilities.
@@ 3 AbilityLens
Table 2: The data component detail of AblilityLens. Data of AbilityLens is sourced from 11 public benchmarks. We reduce the sample size in large datasets, where $\rightarrow$ indicates the number of samples before and after reduction. We also provide further decomposition in the supplementary material.
In this section, we elaborate on the details of our AbilityLens. Our AbilityLens includes (1) dataset collection and metrics, and (2) evaluation framework with online and offline modes.
@@ 3.1 Dataset Collection and Metrics
Dataset collection. A major challenge lying in building a comprehensive ability-focused evaluation is the lack of a unified evaluation protocol, as model and benchmark creators often implement custom pipelines that vary widely in data preparation, output postprocessing, and metric calculation. Therefore, we develop our AbilityLens based on LMMs-Eval [43], a comprehensive evaluation suite for VLMs that supports over 90 VLM benchmarks. In detail, as illustrated in Table 2, AbilityLens focuses on 6 core perception abilities following prior works [12, 40]: counting, OCR, attribute recognition, entity extraction, grounding, and structural data understanding. We select relevant 11 datasets from LMMs-Eval to construct a uniform one for each ability. The included benchmarks encompass MME [12], MMBench [40], SeedBench [18], MuirBench [37], Mirb [45], MMStar [9], OCRBench [24], SynthDog [17], ChartQA [28], AI2D [16], and ConBench [44], covering diverse domain, question types and metrics. To achieve efficient yet broad-ranging evaluation, we follow LMMs-Eval [43] to remove the redundant samples, allowing users to obtain the results of a 7B MLLM in 0.5 hour with 8 H100 GPUs.
@@
Evaluation metric. For each ability, we define two performance metrics—model accuracy and model stability—each computed from sub-metrics extracted from the source benchmarks. Because different question types have varying baseline performance levels, representing the expected accuracy from random guessing, we first adjust each sub-metric to account for these differences. For instance, True/False questions have a baseline accuracy of 50% due to the binary choice, while multiple-choice questions may have a lower baseline depending on the number of options. To ensure that the true ability score remains independent of question type, we apply baseline correction to each sub-metric, aligning them to a consistent standard. Let $m_{ij}$ represent the result of model $i$ on the $j$-th sub-metric of an ability, with $\text{UB}_{j}$ denoting the upper bound of the metric and $\text{BL}_{j}$ as its baseline performance [^1] The baseline performance for the VQA benchmark is set to 0.. The baseline correction for the result $m_{ij}$ is then defined as:
$\tilde{m}_{ij}=\frac{m_{ij}-\text{BL}_{j}}{\text{UB}_{j}-\text{BL}_{j}}.$ (1)
@@
The accuracy of model $i$ for each ability, denoted as $\mathcal{A}_{i}$, is then calculated as a weighted sum of corresponding corrected sub-metrics $\tilde{m}_{ij}$, weighted by their sample counts:
$\mathcal{A}_{i}=\sum_{j}\frac{n_{j}\cdot\tilde{m}_{ij}}{N},$ (2)
@@
where $N=\sum_{j}{n_{j}}$ is the total sample count and $n_{j}$ is the sample count for the $j$-th metric. To assess the model stability of each ability, we first use z-score $z_{ij}$ to measure the relative performance of model $i$ with other model candidates in each sub-metric:
$z_{ij}=\frac{m_{ij}-\mu_{j}}{\sigma_{j}},$ (3)
where $\mu_{j}$ and $\sigma_{j}$ are the mean and variance of the results of model candidates on this sub-metric, respectively. In this paper, we use 18 MLLMs from different series, sizes, sources, and LLMs as the candidates, capturing a wider spectrum of architectural variations and performance characteristics. Then, we measure the standard deviation of the resulting z-scores as the stability score $\mathcal{I}$ of an ability, with higher values indicating greater instability:
$\mathcal{I}_{i}=\operatorname{std}(\mathbf{z}_{i}).$ (4)
@@ 3.2 Evaluation Framework
Figure 2: Monitoring training dynamics with AbilityLens. Unlike traditional methods that focus solely on average scores, often obscuring underlying conflicts, AbilityLens tracks optimization trends for each specific ability, revealing an ability conflict indicated by a significant performance degradation in accuracy and stability of OCR performance after 10,000 steps.
@@
Offline Evaluation Mode provides a comprehensive assessment of each model’s perception capabilities in accuracy and stability, as shown in Table 3 and 4, respectively. This approach effectively reduces evaluation bias and variance caused by the single benchmark, enabling a more robust and holistic comparison across models. We evaluate state-of-the-art MLLMs with various model sizes and training data, including both commercial and open-source models, across six core abilities.
Online Evaluation Mode enables real-time monitoring of training dynamics throughout model development, leveraging AbilityLens for a more unified and comprehensive assessment compared to existing benchmarks. By evaluating checkpoints at various training stages, we can (1) assess the impact of training configuration adjustments, (2) identify optimal checkpoints for specific abilities, (3) track trends in ability-specific optimization, and (4) detect potential ability conflicts where some abilities meet performance degradation in both accuracy and stability after further training. Addressing these conflicts during training can lead to more balanced and stable multimodal capabilities, ultimately enhancing the overall effectiveness of the models. In Figure 2, we illustrate AbilityLens in action through a reimplementation of LLaVA-OV-SI’s [19] single-stage training process. Traditional approaches, represented by the fourth subfigure in Figure 2, focus solely on average scores and often obscure underlying conflicts. In contrast, AbilityLens offers a capability-decomposed analysis tool that enables researchers to gain deeper insights and make targeted improvements in their VLMs. For example, in the second figure of Figure 2, after 10,000 training steps, the OCR ability displays a marked performance degradation in both accuracy and stability, revealing an ability conflict that conventional methods might overlook. We will discuss using existing fine-tuning and model merging methods to solve this issue in Sec 5.
Table 3: Accuracy comparison. Accuracy $\uparrow$ of 18 MLLMs on AbilityLens where cells highlighted in green indicate the best performance, and red indicate the worst performance.
@@
Qwen2VL-2b 48.28 73.36 35.00 82.39 56.61 46.36 57.00
Qwen2VL-7b 50.95 79.29 47.12 85.10 62.66 67.56 65.45
Qwen2VL-72b 55.84 86.09 58.70 86.43 65.11 79.32 71.92
Qwen2.5VL-3b 50.27 77.49 40.01 82.56 60.70 56.10 61.19
Qwen2.5VL-7b 50.52 80.10 47.71 86.19 62.60 72.44 66.59
Qwen2.5VL-72b 56.53 85.70 59.62 86.82 67.11 79.20 72.49
@@
Table 4: Stability comparison. Stability score $\downarrow$ of 18 MLLMs on AbilityLens, with cells highlighted in red for the highest and green for the lowest scores. Lower scores indicate better stability.
@@
LLaVA1.5-7b 72.54 128.40 42.32 110.40 85.30 46.73 80.95
LLaVA1.6-7b 48.00 72.73 31.47 57.29 28.37 43.95 46.96
LLaVA-OV-0.5b 67.07 110.5 60.61 50.00 82.58 73.21 74.00
LLaVA-OV-7b 65.09 62.05 49.28 26.81 72.56 18.62 59.64
LLaVA-OV-SI-7b 115.80 34.17 86.64 42.46 53.20 25.74 88.54
LLaVA-Video-7b 152.20 60.16 78.28 132.30 28.56 79.74 56.43
LLaVA-OV-72b 96.88 57.49 58.85 40.45 42.29 42.63 56.43
Qwen2VL-2b 47.50 89.28 56.53 20.97 64.92 46.80 54.33
Qwen2VL-7b 44.08 45.69 57.59 12.01 46.24 27.55 38.86
Qwen2VL-72b 88.18 24.78 51.52 31.40 49.34 27.82 44.56
Qwen2.5VL-3b 45.16 72.17 60.24 18.88 52.40 47.09 49.32
Qwen2.5VL-7b 45.10 46.70 56.91 14.09 47.28 22.17 38.71
Qwen2.5VL-72b 82.81 29.70 49.62 39.89 43.10 28.33 45.57
InternVL2-8b 44.16 17.34 47.93 9.74 54.60 60.48 38.54
InternVL3-8b 49.12 24.63 44.41 13.82 51.17 62.05 40.87
gpt-4-vision-preview 85.72 77.08 93.53 80.34 52.74 55.78 64.20
gpt-4o-2024-08-06 86.98 24.01 57.06 8.30 12.29 36.48 37.52
claude-3-5-sonnet-2024102 84.57 24.78 59.68 10.85 25.86 27.80 38.92